Diffusion Morning Report 02/01/05

Ted Berg tberg at snf.stanford.edu
Tue Feb 1 06:49:15 PST 2005


Good Morning All,
    Looking pretty good this morning.  GeH4 and HCl for epi cylinders 
changed yesterday. Leak on Epi repaired.  SiGe- SiH4 zero drift.
Ebeam still in High Vac.  Everything else looks OK for now.



. 	Date: 	2/1/2005 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing 	 
Tylan 8-SiGe 	yellow 	SiH4 leak 	12.33.0/30
Tylan 9-Poly 	Up 	  	17.59/30
Tylan 10-Nitride 	Up 	  	11.03/30
Tylan 11-BPSG-LTO 	Up 	  	15.93/30
Tylan 12-FGA 	Up 	  	 
Tystar 	yellow 	  	3.48/30
Doped Poly 	GONE 	  	 
TEOS 	up 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 480 PSI
  	  	  	 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050201/7fa13784/attachment.html>


More information about the diffusion mailing list